Find an equation for the plane perpendicular to the vector A=2i+3j+6k and passing through the terminal point of the vector B=1+5j+3k
1
Expert's answer
2019-10-16T13:31:28-0400

The equation of the plane with normal "\bold n = (A, B, C)" that goes through point "\bold r_0 = (x_0, y_0, z_0)" is "\bold n \cdot (\bold r - \bold r_0) = A (x- x_0) + B(y - y_0) + C (z- z_0) = 0".

In our case, vector "\bold A = (2, 3, 6)" can be used as normal vector, and the terminal point of vector "\bold B" can be used as "\bold r_0". Hence, the equation of the plane is "2 (x-1) + 3 (y-5) + 6 (z - 3) = 0", or simplified "2 x + 3 y + 6z - 35 = 0".